The mean grade point average for statistical inference I course in a particular session is 2.45 with variance of 0.4761 it is believed that statistics majors have higher grade point average than the overall mean. A sample of 100 statistics major had a mean grade point average of 2.65. Test his claim at 1% level of significance


1
Expert's answer
2022-02-01T11:26:12-0500

"H_0:\mu=2.45" , majors have not higher grade point average than the overall mean

"H_a:\mu>2.45" , majors have higher grade point average than the overall mean


"z=\frac{\overline{x}-\mu}{\sigma/\sqrt{n}}=\frac{2.65-2.45}{\sqrt{0.4761}/10}=2.90"


critical value:

"z_{crit}=2.576"


Since "z>z_{crit}" , we have to reject the null hypothesis. Majors have higher grade point average than the overall mean.